OpenStudy (cooltowl):

You heated 2.92g of baking soda in a crucible to produce 1.810g of a solid product. The product could be NaOH, Na2CO3, or Na2O. What mass of NaOH could you have produced?

OpenStudy (aaronq):

the equation for the decomposition of sodium bicarbonate into sodium hydroxide also produces carbon dioxide. Balanced, the equation is: \(\sf NaHCO_3\rightarrow NaOH+CO_2\) Lets work out the theoretical yield with the mass of NaHCO3 1. First convert to moles with: \(\sf moles=\dfrac{mass}{Molar~mass}\) 2. Next use a ratio of the moles of both chemical species and their stoichiometric coefficients: \(\sf \dfrac{moles~of~NaHCO_3}{NaHCO_3's~coefficient}=\dfrac{moles~of~NaOH}{NaOH's~coefficient}\) Plug in moles of NaHCO3, the coefficients and solve for moles of NaOH (because the coefficients are both 1 in this example, we can omit this^ last step) 3. Lastly, convert to mass using equation from 1, solve for mass of NaOH this time.